Consider a portfolio that is 60% (wS) in the S&P 500 index and 40% (wB) in the JPM bond index.

Suppose an analyst estimates that E[rS] = 15%, σs = 20%, E[rB] = 10%, σB =12%, and ρ = 0%.

1) Calculate the portfolio's standard deviation given that the correlation is zero:

2) Calculate the weighted average of the standard deviations of the S&P 500 index and the JPM bond index:
Answer #1

1]

standard deviation for a two-asset portfolio σp = (w12σ12 + w22σ22 + 2w1w2Cov1,2)1/2

w1 = weight of Asset 1

w2 = weight of Asset 2

σ12 = variance of Asset 1

σ22 = variance of Asset 2

Cov1,2 = covariance of returns between Asset 1 and Asset 2

Cov1,2 = ρ1,2 * σ1 * σ2, where ρ1,2 = correlation of returns between Asset 1 and Asset 2

standard deviation = ((0.60)2(0.20)2 + (0.40)2(0.12)2 + (2)(0.60)(0.40)(0)(0.20)(0.12))1/2

standard deviation = 12.92%

2]

weighted average of the standard deviations = (0.60 * 0.20) + (0.40 * 0.12) = 16.80%